OpenStudy (anonymous):

I would really appreciate it if someone would help me with this question. Tina wants to save money for school. Tina invests $1,000 in an account that pays an interest rate of 8% How many years will it take for the account to reach $1,300? Round your answer to the nearest hundredth. Using the formula: A=P(1+r)^t

OpenStudy (anonymous):

@Hoslos

OpenStudy (anonymous):

@anelgomez__ what do you get using the aforementioned formula?

OpenStudy (anonymous):

the answer that i had gotten was wrong @Hoslos

OpenStudy (anonymous):

Did you do\[\ln A=t \ln P(1+r)\rightarrow t =\frac{ \ln A }{ \ln P(1+r) }\]? Which subsequently gives \[\frac{ \ln 1300 }{ \ln 1000(1+0.08) }=?\] If so or if you understood, find '?'

OpenStudy (anonymous):

is it 1.12 @Hoslos

OpenStudy (anonymous):

It is ln1300/(ln1000)*1.08=0.96 @anelgomez__

OpenStudy (anonymous):

Thank you!!(: @Hoslos

OpenStudy (anonymous):

No problem @anelgomez__.